A particle with a charge of q = -5.00 nC is moving in a uniform magnetic field of B⃗ =( -1.27 T ) z^. The magnetic force on the particle is measured to be F⃗ =( −7.60×10−7 N )y^. Calculate vx, the x component of the velocity of the particle.
Answer #1

here ,

q = -5 nC

B = -1.27 T z

F = -7.6 *10^-7 y

for the force

F = q * v X B

-7.6 *10^-7 = -5 *10^-9 * (v * 1.27)

solving for v

vx = 119.6 m/s

the velocity of the particle is 119.6 m/s